# The stability of propagating plane inertial waves in rotating fluids

### v1

@@ II.2 Floquet Solutions
Solutions of the Floquet system in Equations (II.1)-(18) are examined below using a numerical solver. The numerical solution to the infinite set of coupled equations is obtained by truncating the sum over $n$ with a finite number of modes $N$ in the range $N/2<n\leq N/2$. The system is further reduced to $3N$ equations in the variables $u_{n},v_{n},w_{n}$ by eliminating the pressure $p_{n}$ using the incompressibility constraint in Equation (18). The resulting system is a sparse $3N\times 3N$ matrix whose eigenvalues and eigenvectors can be found using standard numerical software. The number of modes $N$ is chosen to be large enough that the eigenvector components are negligible near the truncation mode $|n|\sim N/2$. The Ekman number $\mathrm{Ek}$ is chosen to be sufficiently low such that the viscous damping rate at all horizontal wavenumbers is much smaller than the typical instability growth rate $\nu(\alpha^{2}+\beta^{2})\ll\sigma\sim A^{\prime}|\omega|$, which can be written as the condition $\mathrm{Ek}\ll A^{\prime}k^{2}/(\alpha^{2}+\beta^{2})$. We focus on how the amplitude and frequency of the primary IW affects the growth rate and wavenumber orientation of the most unstable perturbations.
@@
Figure 1: The maximum growth rate $\sigma^{\max}$ of unstable perturbations predicted by Floquet theory as a function of the perturbation’s horizontal wavenumbers $\alpha,\beta$ for primary waves with frequencies $\omega/2\Omega=0.17,0.71,0.98$ (a-c), corresponding to angles $\theta=80^{\circ},45^{\circ},10^{\circ}$. The growth rate is normalized by $A^{\prime}|\omega|$ and the Floquet parameter is fixed to $\gamma=0$. All primary waves have the same normalized amplitude $A^{\prime}=0.3$ and helicity $s=1$.
Figure 2: The maximum normalized growth rate of unstable perturbations for primary waves with amplitudes $A^{\prime}=0.03,0.3,3.0$ (a-c), all with the same frequency $\omega=\Omega$ (angle $\theta=60^{\circ}$) and helicity $s=1$. The Floquet parameter is $\gamma=0$.
@@
We first examine the effect of varying the dimensionless wave frequency $\omega/2\Omega$, or equivalently the angle $\theta$ between $\bm{k}$ and $\bm{\Omega}$. Figure 1 shows Floquet stability diagrams for primary waves with frequencies $\omega/2\Omega=0.17,0.71,0.98$, (angles $\theta=80^{\circ},45^{\circ},10^{\circ}$), all with the same amplitude $A^{\prime}=0.3$ and helicity $s=1$. The Floquet diagrams show the maximum growth rate $\sigma^{\max}$ obtained at each horizontal wavenumber $\alpha,\beta$ using the numerical Floquet solver with $N=128$ modes. The growth rate is normalized by $A^{\prime}|\omega|$ and the Floquet parameter is fixed to $\gamma=0$. A characteristic feature of these plots is the ribbon patterns of the most unstable horizontal wavenumbers. The center of each ribbon corresponds to an exact resonance of the PSI at small amplitudes $A^{\prime}\ll 1$, which is confirmed by an asymptotic analysis in Appendix B (see Fig. 10). One can see that the orientations of the most unstable horizontal wavenumbers has a strong dependence on the primary wave frequency. Lower frequency waves ($\bm{k}$ nearly orthogonal to $\bm{\Omega}$) are more unstable to perturbations with wavevectors in the $\hat{\bm{y}}$ direction (with $\alpha\sim 0$), orthogonal to the $\bm{k}-\bm{\Omega}$ plane. In contrast, higher frequency waves ($\bm{k}$ nearly parallel to $\bm{\Omega}$) are unstable to perturbations whose wavenumbers orientations are more isotopically distributed in the plane perpendicular to $\bm{k}$. This is in agreement with the linear stability analysis in the short-wavelength limit in [39, see also Appendix B]. While these results are shown only for $\gamma=0$, we have found that $\gamma\neq 0$ slightly changes the locations of the ribbons but does not change the observed trends with $\omega$ and $A^{\prime}$.
@@
We note a few additional properties of the Floquet diagrams that we have observed numerically. The diagrams all appear symmetric across the $\beta=0$ axis, despite no obvious symmetry for $\alpha\rightarrow-\alpha$ in the Floquet system II.118. We also find that the most unstable perturbations always have $\alpha=0$ and that the local maxima in $\sigma^{\max}/A^{\prime}|\omega|$ on the $\beta$-axis are independent of the wave frequency, which we confirmed with extra numerical scans. The main effect of increasing the wave frequency is thus the increase in the growth rates of perturbations with finite $\alpha$, resulting in more isotropic ribbon patterns. Lastly, we note that the Floquet diagrams are unchanged when flipping the primary wave helicity $s\rightarrow-s$.
@@
Next, we vary the primary wave amplitude $A^{\prime}$. Figure 2 shows Floquet stability diagrams for primary waves with amplitudes $A^{\prime}=0.03,0.3,3.0$, all with the same frequency $\omega=\Omega$ (angle $\theta=60^{\circ}$) and helicity $s=1$. The main effect of moderately increasing the wave amplitude is the increased widths of the ribbons, as one can see by comparing panels (a) and (b) with $A^{\prime}=0.03$ and $A^{\prime}=0.3$, respectively. However, the predictions of the PSI break down at large amplitudes $A^{\prime}\sim 1$. This is evident in the case with $A^{\prime}=3.0$ in panel (c), where the observed patterns no longer match the ribbons seen at low $A^{\prime}$ and the fastest growing perturbations have horizontal wavenumbers $\alpha=0$, $\beta\approx 1$ that are comparable to those of the primary wave.
@@
Figure 3: The maximum normalized growth rate $\sigma^{\max}/A^{\prime}|\omega|$ versus the primary wave frequency $\omega/2\Omega$ for a range of amplitudes $0.1\leq A^{\prime}\leq 3.0$. For each $\omega$ and $A^{\prime}$, $\sigma^{\max}$ is determined by finding the fastest growth rate among perturbations with wavenumbers in the range $16\leq\alpha\leq 16$, $0\leq\beta\leq 16$, and $\gamma=0$ using the numerical Floquet solver. The magnitude of $\sigma^{\max}/A^{\prime}|\omega|$ is nearly independent of $\omega$ and is weakly dependent on $A^{\prime}$. At low $A^{\prime}\ll 1$, the growth rate based on Floquet theory agrees with the maximum growth rate of the PSI, $\sigma_{\rm PSI}/A^{\prime}|\omega|=3\sqrt{3}/16$, marked by the dashed, gray line.
@@
An interesting feature of the Floquet solutions is that the maximum value of the normalized growth rate $\sigma^{\max}/A^{\prime}|\omega|$ across all $\alpha,\beta$ is nearly independent of both $A^{\prime}$ and $\omega/2\Omega$, even for $A^{\prime}\gtrsim 1$. It robustly has a magnitude of $\sim 0.3$. A broader scan of the parameter space $(A^{\prime},\omega/2\Omega)$ confirms this feature. Figure 3 shows the maximum normalized growth rate across a range of primary wave amplitudes $0.1\leq A^{\prime}\leq 3$ and frequencies $0.1\leq\omega/2\Omega\leq 0.9$. At low wave amplitudes, its nearly constant value matches the maximum PSI growth rate $\sigma_{\rm PSI}/A^{\prime}|\omega|=3\sqrt{3}/16\approx 0.325$ [39, see also Appendix B]. At high wave amplitudes, the maximum normalized growth rate deviates remarkably little from the PSI prediction (derived for $A^{\prime}\ll 1$), decreasing only slightly to $\approx 0.28$ for wave amplitudes of $A^{\prime}\sim 3$. For even larger $A^{\prime}\gtrsim 3$, the instability growth rate exceeds the primary wave frequency $\sigma^{\max}>\omega$, implying that propagating IWs with such large amplitudes would be difficult to observe in any application.
@@ III Simulations of Plane Inertial Waves
Investigating the nonlinear saturation of the instabilities of a propagating IW and the subsequent turbulent evolution of the flow requires using DNS. We carry out DNS in both the comoving and the rotating frames, leveraging the numerical advantages of each frame. Simulations in the comoving frame enable a cross-validation between the DNS and the Floquet solvers. The discrete horizontal wavenumbers in the comoving frame of the DNS correspond exactly to the $\alpha,\beta$ used in the Floquet analysis, allowing a direct verification of the Floquet theory. Simulations in the rotating frame are better suited to follow the nonlinear evolution of the flow. The rotating frame has the rotation vector exactly aligned with the numerical grid, allowing geostrophic modes with $k_{z}=0$ be exactly captured. Our focus will be to examine how the energy of the primary wave is partitioned into the geostrophic and the non-geostrophic components of the turbulent flow as the wave breaks down.
@@
The Navier-Stokes equation for the velocity field is solved as an initial value problem using the Dedalus code [45]. The domain is a cube of side length $L=16\pi$ with periodic boundary conditions, and the flow is represented as a truncated sum of Fourier modes $\bm{u}=\sum_{\bm{k}}\bm{u}_{\bm{k}}(t)\exp(i\bm{k}\cdot\bm{x})$. Time stepping is carried out with a third-order, implicit-explicit Runge-Kutta method [46] where linear terms are treated implicitly, non-linear terms are treated explicitly, and the timestep $dt$ is set as a small fraction of the shortest growth timescale, $dt\ll 1/\sigma^{\max}$. The nonlinear terms of the Navier-Stokes equations are evaluated in real space with the standard $2/3$ dealiasing rule. All runs use a spectral resolution of $N=256^{3}$ Fourier modes.
@@
For simulations in the comoving frame, the time-independent base flow $\bm{U}^{\prime}(\phi)$ is initialized following Equation (7), which fits $L/2\pi=8$ wavelengths along the vertical $\hat{\phi}$ direction. The wave frequency, $\omega/2\Omega=s\cos\theta$, is controlled by the tilt of the rotation vector $\bm{\Omega}^{\prime}=(-\sin\theta\hat{\bm{x}}+\cos\theta\hat{\bm{\phi}})/|\cos\theta|$. For simulations in the rotating frame, the base flow $\bm{U}(\bm{x},t=0)$ is initialized following Equation (2). This excites an IW with a time-dependent flow field $\bm{U}(\bm{x},t)$ propagating inside the domain. The wave frequency is controlled by the choice of wavenumber $\bm{k}=(k_{x},0,k_{z})$ in the xxzz plane, where only certain frequencies $\omega/2\Omega=sk_{z}/k$ (here $k=\sqrt{k_{x}^{2}+k_{z}^{2}}$) can be chosen based on the discrete Fourier grid. We choose wavenumbers such that $kL/2\pi\sim 8$ to achieve a similar scale separation between the wavelength and the box scale as in the comoving frame. In both setups, the Ekman number is chosen to damp unstable modes at the largest wavenumbers in the box $k_{\max}=\pi N/L$ by setting $\mathrm{Ek}=CA^{\prime}/k_{\max}^{2}$, where $C\sim 1$ is an adjustable “safety” factor used to control the level of damping. The helicity of all primary waves studied below is $s=1$.
@@ III.1 Verification of Linear Theory with DNS in the Comoving Frame
We first cross-validate the DNS in the comoving frame and the Floquet solver by comparing the growth rates obtained independently by the two methods. The growth rates in the DNS are measured by fitting the exponential growth of the energy in the discrete horizontal Fourier modes, which correspond to the horizontal wavenumbers ($\alpha,\beta$) in the Floquet analysis after they have been rescaled by a multiplicative factor of $8$, equal to ratio of the box scale to the wavelength of the primary wave. The energy $E_{\alpha,\beta}(t)$ at each $(\alpha,\beta)$ is computed by summing over all vertical wavenumbers $k_{z}$, that is $E_{\alpha,\beta}(t)=0.5\sum_{k_{z}}|\bm{u}^{\prime}_{\bm{k}}(\alpha,\beta,k_{z},t)|^{2}$. The growth of $E_{\alpha,\beta}(t)$ is dominated by the fastest growing perturbation with the given $(\alpha,\beta)$ and should closely match the value of $2\sigma^{\max}(\alpha,\beta)$ predicted by Floquet theory. The fitting is done over a time interval where the velocity field perturbations grow from an initial amplitude of $|\bm{u}^{\prime}|\sim 10^{-8}A^{\prime}$ to a final amplitude of $|\bm{u}^{\prime}|\sim 10^{-2}A^{\prime}$. The initial velocity field is a Gaussian random noise field in real space whose solenoidal component is projected away.
@@
Figure 4: The DNS in the comoving frame is cross-validated against the Floquet theory by comparing the maximum growth rate $\sigma^{\max}$ in the $\alpha-\beta$ plane for a primary wave with amplitude $A^{\prime}=0.3$ and frequency $\omega/2\Omega=0.71$ (angle $\theta=45^{\circ}$). Panel (a) shows the Floquet prediction for $\sigma^{\max}$, where the fastest growing perturbation at each $\alpha,\beta$ is found by also scanning all possible $\gamma$ allowed on the discrete Fourier grid of the DNS. Panel (b) shows the $\sigma^{\max}$ measured at each $\alpha,\beta$ in the DNS.
@@
The comparison of $\sigma^{\max}$ in the $\alpha-\beta$ plane is shown in Figure 4 for a primary IW with $A^{\prime}=0.3$, and $\omega/2\Omega=0.71$ (angle $\theta=45^{\circ}$). The figure shows only the zoomed-in region $0\leq\alpha,\beta\leq 2$ for clearer visual inspection. Although the measurement of the growth rates in the DNS can be somewhat noisy (in particular for the slower growing modes), the DNS result and the Floquet predictions are in good agreement. Note that all available $0\leq\gamma<1$ can be excited and the Floquet solver searches over $\gamma=j/8$ with all possible integers $0\leq j<8$ that exist on the numerical Fourier grid (by contrast, Figure 1 and  2 only showed the maximum growth rate among perturbations with $\gamma=0$). We find that the different ribbon structures associated with each $\gamma$ end up almost completely covering the $\alpha-\beta$ plane. Perturbations with nearly every horizontal wavenumber ($\alpha,\beta$) have a significant growth rate, with the exception of the low horizontal wavenumbers around $\sqrt{\alpha^{2}+\beta^{2}}\lesssim 0.5$ whose wavelengths are larger than the wavelength of the primary IW.
@@
Figure 5: 2D energy spectra $E_{\alpha,\beta}$ during the linear phase of the DNS in the comoving frame for three primary waves with different frequencies $\omega/2\Omega$, but the same amplitude $A^{\prime}=0.3$. The fastest growing perturbations gain the most energy and their distribution in the $\alpha-\beta$ plane depends on the wave frequency in a similar manner as the Floquet predictions in Figure 1. The contours of constant energy are shown as the black dashed lines at integer values of $\log_{10}(E_{\alpha,\beta})=i$ in the range $9\leq i\leq-15$.
@@
The Floquet theory prediction for how the primary wave frequency affects the wavenumber orientation of the most unstable perturbations can also be confirmed by examining the energy spectra $E_{\alpha,\beta}$ for the three fiducial frequencies $\omega/2\Omega=0.17,0.71,0.98$, as shown in Figure 5. The energy spectra are computed at similar times during the linear phase of each simulation and the $\alpha,\beta$ with the largest energies corresponds to the perturbations that have grown the fastest from their initial values of $E_{\alpha,\beta}\sim 10^{-16}$. One can see that the energy spectra in Figure 5 qualitatively have the same features and dependence on $\omega/2\Omega$ as the Floquet predictions in Figure 1. The wavenumber orientations of unstable perturbations shift from being preferentially aligned with the $\hat{\bm{y}}^{\prime}$ direction for lower frequency waves to being more isotropically distributed for higher frequency waves. The energy in horizontal wavenumbers larger than $\sqrt{\alpha^{2}+\beta^{2}}\sim 8$ is small because the viscous damping at these large wavenumbers becomes significant for the Ekman number $\mathrm{Ek}\sim 10^{-3}$ needed to resolve these simulations.

### v2

@@ II.2 Floquet Solutions
Figure 1: The maximum normalized growth rate $\sigma^{\max}/A^{\prime}|\omega|$ of unstable perturbations predicted by Floquet theory as a function of the perturbation’s horizontal wave vector components $\alpha,\beta$ for primary waves with frequencies $\omega/2\Omega=0.17,0.71,0.98$ (a-c), corresponding to angles $\theta=80^{\circ},45^{\circ},10^{\circ}$. The Floquet parameter is fixed to $\gamma=0$ and all primary waves have the same normalized amplitude $A^{\prime}=0.3$, helicity $s=1$, and $\mathrm{Ek}=10^{-6}$.
Solutions of the Floquet system in Equations (II.1)-(18) are examined below using a numerical solver. The numerical solution to the infinite set of coupled equations is obtained by truncating the sum over $n$ with a finite number of modes $N$ in the range $N/2<n\leq N/2$. The system is further reduced to $3N$ equations in the variables $u_{n},v_{n},w_{n}$ by eliminating the pressure $p_{n}$ using the incompressibility constraint in Equation (18). The resulting system is a sparse $3N\times 3N$ matrix whose eigenvalues and eigenvectors can be found using standard numerical software. The number of modes $N$ is chosen large enough that eigenvector components are negligible near the truncation mode $|n|\sim N/2$. The Ekman number $\mathrm{Ek}$ is chosen to be sufficiently small such that the viscous damping rate at all horizontal wave vectors is much smaller than the typical instability growth rate $\mathrm{Ek}(\alpha^{2}+\beta^{2})\ll\sigma^{\prime}\sim A^{\prime}$. We focus on how the amplitude and frequency of the primary IW affect the growth rate and wave vector orientation of the most unstable perturbations.
Figure 2: The maximum normalized growth rate of unstable perturbations with Floquet parameters $\gamma=0,1/3,1/4$ (a-c) for the same primary wave with $A^{\prime}=0.3$, $\omega/2\Omega=0.71$, $s=1$, and $\mathrm{Ek}=10^{-6}$. The panels focus on the origin where the effects of $\gamma$ are most pronounced since $\alpha\sim\beta\sim\gamma$. The $\gamma=0$ case in panel (a) is a zoomed-in region of panel (b) in Figure 1.
@@
We first vary the dimensionless wave frequency $\omega/2\Omega$, or equivalently the angle $\theta$ between $\bm{k}$ and $\bm{\Omega}$. Figure 1 shows Floquet stability diagrams for primary waves with frequencies $\omega/2\Omega=0.17,0.71,0.98$, (angles $\theta=80^{\circ},45^{\circ},10^{\circ}$), all with the same amplitude $A^{\prime}=0.3$ and helicity $s=1$. The Floquet diagrams show the maximum normalized growth rate $\sigma^{\max}/A^{\prime}|\omega|$ obtained at each horizontal wave vector with components $\alpha,\beta$ using the numerical Floquet solver with $N=128$ modes and a fixed Floquet parameter $\gamma=0$ (the dimensional growth rate is related to the non-dimensional growth rate by $\sigma=\sigma^{\prime}|\omega|$). A characteristic feature of these plots is the ribbon patterns of the most unstable horizontal wave vectors. The center of each ribbon corresponds to an exact resonance of the PSI at small amplitudes $A^{\prime}\ll 1$, which is confirmed by an asymptotic analysis in Appendix B (see Figure 11). One can see that the orientations of the most unstable horizontal wave vectors have a strong dependence on the primary wave frequency. Lower frequency waves ($\bm{k}$ nearly orthogonal to $\bm{\Omega}$) are more unstable to perturbations with wavevectors in the $\hat{\bm{y}}$ direction (with $\alpha\sim 0$), orthogonal to the $\bm{k}-\bm{\Omega}$ plane. In contrast, higher frequency waves ($\bm{k}$ nearly parallel to $\bm{\Omega}$) are unstable to perturbations whose wave vector orientation is more isotopically distributed in the plane perpendicular to $\bm{k}$. This is in agreement with the linear stability analysis in the short-wavelength limit in [52, see also Appendix B].
While the above results are for $\gamma=0$, a non-zero Floquet parameter does not change the trends observed with $\omega$ and $A^{\prime}$. Physically, $\gamma$ accounts for modes that are not periodic along $\hat{\phi}$ with the primary wave’s wavelength. Since the Floquet parameter is bounded as $0\leq\gamma<1$, varying $\gamma$ slightly shifts the instability ribbons in the $\alpha-\beta$ plane at large wavenumbers ($\alpha^{2}+\beta^{2}\gg\gamma^{2}$); however, it does not affect the maximum value of the growth rate (Appendix B). This can be observed in Floquet diagrams (Figure 2) where $\gamma$ is varied for the $\omega/2\Omega=0.71$ wave.
@@
We note a few additional properties of the Floquet diagrams that we have observed numerically. The diagrams all appear symmetric across the $\beta=0$ axis for $\gamma=0$, despite no obvious symmetry for $\alpha\rightarrow-\alpha$ in the Floquet system II.118. However, this symmetry is broken by non-zero values of the Floquet parameter $\gamma\neq 0$. We also find that the most unstable perturbations always have $\alpha=0$ and that the local maxima in $\sigma^{\max}/A^{\prime}|\omega|$ on the $\beta$-axis are independent of the wave frequency, which we confirmed with extra numerical scans. The main effect of increasing the wave frequency is thus the increase in the growth rates of perturbations with finite $\alpha$, resulting in more isotropic ribbon patterns. Lastly, flipping the primary wave helicity $s\rightarrow-s$ does not change the Floquet diagrams.
@@
Next, we vary the primary wave amplitude $A^{\prime}$. Figure 3 shows Floquet stability diagrams for primary waves with amplitudes $A^{\prime}=0.03,0.3,3.0$, all with the same frequency $\omega=\Omega$ (angle $\theta=60^{\circ}$) and helicity $s=1$. The main effect of moderately increasing the wave amplitude is the increased widths of the ribbons, as one can see by comparing panels (a) and (b) with $A^{\prime}=0.03$ and $A^{\prime}=0.3$, respectively. However, the predictions of the PSI break down at large amplitudes $A^{\prime}\sim 1$. This is evident in the case with $A^{\prime}=3.0$ in panel (c), where the observed patterns no longer match the ribbons seen at low $A^{\prime}$ and the fastest growing perturbations have horizontal wave vectors $\alpha=0$, $\beta\approx 1$ that are comparable to those of the primary wave.
Figure 3: The maximum normalized growth rate of unstable perturbations for primary waves with amplitudes $A^{\prime}=0.03,0.3,3.0$ (a-c), all with the same frequency $\omega=\Omega$ (angle $\theta=60^{\circ}$), helicity $s=1$, and $\mathrm{Ek}=10^{-6}$. The Floquet parameter is $\gamma=0$.
@@
Figure 4: The maximum normalized growth rate $\sigma^{\max}/A^{\prime}|\omega|$ versus the primary wave frequency $\omega/2\Omega$ for a range of amplitudes $0.1\leq A^{\prime}\leq 3.0$. For each $\omega$ and $A^{\prime}$, the $\sigma^{\max}$ is determined by finding the fastest growth rate among perturbations with wave vector components in the range $16\leq\alpha\leq 16$, $0\leq\beta\leq 16$, and $\gamma=0$ using the numerical Floquet solver. The magnitude of $\sigma^{\max}/A^{\prime}|\omega|$ is nearly independent of $\omega$ and is weakly dependent on $A^{\prime}$. At low $A^{\prime}\ll 1$, the growth rate based on Floquet theory agrees with the maximum growth rate of the PSI, $\sigma_{\rm PSI}/A^{\prime}|\omega|=3\sqrt{3}/16$, marked by the dashed, gray line.
@@
An interesting feature of the Floquet solutions is that the maximum normalized growth rate $\sigma^{\max}/A^{\prime}|\omega|$ across all $\alpha,\beta$ is nearly independent of both $A^{\prime}$ and $\omega/2\Omega$, even for $A^{\prime}\gtrsim 1$. It robustly has a magnitude of $\sim 0.3$. A broader scan of the parameter space $(A^{\prime},\omega/2\Omega)$ confirms this feature. Figure 4 shows the maximum normalized growth rate across a range of primary wave amplitudes $0.1\leq A^{\prime}\leq 3$ and frequencies $0.1\leq\omega/2\Omega\leq 0.9$. At low wave amplitudes, its nearly constant value matches the maximum PSI growth rate $\sigma_{\rm PSI}/A^{\prime}|\omega|=3\sqrt{3}/16\approx 0.325$ [52, see also Appendix B]. At high wave amplitudes, the maximum normalized growth rate deviates remarkably little from the PSI prediction (derived for $A^{\prime}\ll 1$), decreasing only slightly to $\approx 0.28$ for wave amplitudes of $A^{\prime}\sim 3$. This is in agreement with the analysis of [1] where the growth rate of short-wavelength perturbations was also found to begin deviating from the PSI prediction for $A^{\prime}\gtrsim 1$. For even larger $A^{\prime}\gtrsim 3$, the instability growth rate exceeds the primary wave frequency $\sigma^{\max}>|\omega|$, implying that propagating IWs with such large amplitudes would be difficult to observe in any application.
@@ III Simulations of Plane Inertial Waves
Investigating the nonlinear saturation of the instabilities of a propagating IW and the subsequent turbulent evolution of the flow requires using DNS. We carry out DNS in both the comoving and the rotating frames, leveraging the numerical advantages of each frame. Simulations in the comoving frame enable a cross-validation between the DNS and the Floquet solvers. The discrete horizontal wave vector components in the comoving frame of the DNS correspond exactly to the $\alpha,\beta$ used in the Floquet analysis, allowing a direct verification of the Floquet theory. Simulations in the rotating frame are better suited to follow the nonlinear evolution of the flow. The rotating frame has the rotation vector exactly aligned with the numerical grid, allowing geostrophic modes with $k_{z}=0$ to be exactly captured. Our focus is to examine how the energy of the primary wave is partitioned into the geostrophic and the non-geostrophic components of the turbulent flow as the wave breaks down.
@@
The Navier-Stokes equation for the velocity field is solved as an initial value problem using the Dedalus code [22]. The domain is a cube of side length $L=16\pi$ with periodic boundary conditions, and the flow is represented as a truncated sum of Fourier modes $\bm{u}=\sum_{\bm{k}}\bm{u}_{\bm{k}}(t)\exp(\mathrm{i}\bm{k}\cdot\bm{x})$. Time stepping is carried out with a third-order, implicit-explicit Runge-Kutta method [4] where linear terms are treated implicitly, non-linear terms are treated explicitly, and the timestep $dt$ is set as a small fraction of the shortest growth timescale, $dt\ll 1/\sigma^{\max}$. The nonlinear terms are evaluated in real space with the standard $2/3$ dealiasing rule. All runs use a resolution of $N=256^{3}$ Fourier modes.
@@
For simulations in the comoving frame, the time-independent base flow $\bm{U}^{\prime}(\phi)$ is initialized following Equation (7), which fits $L/2\pi=8$ wavelengths along the vertical $\hat{\bm{\phi}}$ direction. The wave frequency, $\omega/2\Omega=s\cos\theta$, is controlled by the tilt of the rotation vector $\bm{\Omega}^{\prime}=-|\tan\theta|\hat{\bm{x}}+(\cos\theta/|\cos\theta|)\hat{\bm{\phi}}$. For simulations in the rotating frame, the base flow $\bm{U}(\bm{x},t=0)$ is initialized following Equation (2). This results in an IW with a time-dependent flow field $\bm{U}(\bm{x},t)$ propagating inside the domain. The wave frequency is controlled by the choice of wave vector $\bm{k}=(k_{x},0,k_{z})$ in the xxzz plane, where only certain frequencies $\omega/2\Omega=sk_{z}/k$ (here $k=\sqrt{k_{x}^{2}+k_{z}^{2}}$) can be chosen based on the discrete Fourier grid. We choose wavenumbers such that $kL/2\pi\sim 8$ to achieve a similar scale separation between the wavelength and the box scale as in the comoving frame. To seed unstable perturbations of the base flow, the velocity field in both setups is additionally initialized with a Gaussian random noise vector field whose solenoidal components are projected away. The Ekman number is chosen to damp unstable modes at the largest wavenumbers in the box $k_{\max}=\pi N/L$ by setting $\mathrm{Ek}=CA^{\prime}/k_{\max}^{2}$, where $C\sim 1$ is an adjustable “safety” factor used to control the level of damping. Lastly, the helicity of all primary waves studied below is $s=1$.
@@ III.1 Verification of Linear Theory with DNS in the Comoving Frame
We first cross-validate the DNS in the comoving frame and the Floquet solver by comparing the growth rates obtained independently by the two methods. The growth rates in the DNS are measured by fitting the exponential growth of the energy in the discrete horizontal Fourier modes, which correspond to the horizontal wave vectors ($\alpha,\beta$) in the Floquet analysis after they have been rescaled by a multiplicative factor of $8$, equal to the ratio of the box scale to the wavelength of the primary wave. The energy $E_{\alpha,\beta}(t)$ at each $(\alpha,\beta)$ is computed by summing over all vertical wavenumbers $k_{z}$, that is $E_{\alpha,\beta}(t)=0.5\sum_{k_{z}}|\bm{u}^{\prime}_{\bm{k}}(\alpha,\beta,k_{z},t)|^{2}$. The growth of $E_{\alpha,\beta}(t)$ is dominated by the fastest growing perturbation with the given $(\alpha,\beta)$ and should closely match the value of $2\sigma^{\max}(\alpha,\beta)$ predicted by Floquet theory. The fitting is done over a time duration $\Delta t\approx 5/\sigma^{\max}$ at the end of the simulation when the instantaneous growth rates of individual modes become steady and approach their maximum values. This avoids the initial transient when slower growing modes can significantly contribute and adversely affect the comparison with theory.
@@
Figure 5: The DNS in the comoving frame is cross-validated against the Floquet theory for a primary wave with amplitude $A^{\prime}=0.3$ and frequency $\omega/2\Omega=0.71$ (angle $\theta=45^{\circ}$). Panels (a) and (b) show the Floquet prediction for $\sigma^{\max}$, where the fastest growing perturbation at each $\alpha,\beta$ is found for fixed $\gamma=0$ in panel (a) and for all possible $\gamma$ allowed on the discrete Fourier grid of the DNS in panel (b). Panel (c) shows the $\sigma^{\max}$ measured at each $\alpha,\beta$ in the DNS.
@@
The comparison of $\sigma^{\max}$ in the $\alpha-\beta$ plane is shown in Figure 5 for a primary IW with $A^{\prime}=0.3$, and $\omega/2\Omega=0.71$ (angle $\theta=45^{\circ}$). The figure shows only the zoomed-in region $0\leq\alpha,\beta\leq 2$ for clearer visual inspection. Although the measurement of the growth rates in the DNS can be somewhat noisy (in particular for the slower growing modes), the DNS result (panel c) and the Floquet predictions (panel b) are in good agreement. Note that all available $0\leq\gamma<1$ can be excited and the Floquet solver in panel (b) searches over $\gamma=j/8$ with all possible integers $0\leq j<8$ that exist on the numerical Fourier grid. By contrast, Figures 1,  3 and panel (a) of Figure 5 only show the maximum growth rate among perturbations with $\gamma=0$. The combined instability ribbons for varying $\gamma$ almost completely cover the $\alpha-\beta$ plane, resulting in a non-zero growth rate at nearly every horizontal wave vector for the $A^{\prime}$ adopted. The exception is only for low horizontal wavenumbers $\sqrt{\alpha^{2}+\beta^{2}}\lesssim 0.5$ (wavelengths exceeding that of the primary IW).
@@
Figure 6: 2D energy spectra $E_{\alpha,\beta}$ during the linear phase of the DNS in the comoving frame for three primary waves with different frequencies $\omega/2\Omega$, but the same amplitude $A^{\prime}=0.3$. The fastest growing perturbations gain the most energy, and their distribution in the $\alpha-\beta$ plane depends on the wave frequency in a similar manner as the Floquet predictions in Figure 1. The contours of constant energy are shown as the black dashed lines at integer values of $\log_{10}(E_{\alpha,\beta})=i$ in the range $15\leq i\leq-9$.
@@
The Floquet theory prediction for how the primary wave frequency affects the wave vector orientation of the most unstable perturbations can also be confirmed by examining the energy spectra $E_{\alpha,\beta}$ for the three fiducial frequencies $\omega/2\Omega=0.17,0.71,0.98$, as shown in Figure 6. The energy spectra are computed at similar times during the linear phase of each simulation and the $\alpha,\beta$ with the largest energies correspond to the perturbations that have grown the fastest from their initial values of $E_{\alpha,\beta}\sim 10^{-16}$. One can see that the energy spectra in Figure 6 qualitatively have the same features and dependence on $\omega/2\Omega$ as the Floquet predictions in Figure 1. The wave vector orientations of unstable perturbations shift from being preferentially aligned with the $\hat{\bm{y}}^{\prime}$ direction for lower frequency waves to being more isotropically distributed for higher frequency waves. The energy in horizontal wave vectors with wavenumber larger than $\sqrt{\alpha^{2}+\beta^{2}}\sim 8$ is small because the viscous damping at these large wavenumbers becomes significant for the Ekman number $\mathrm{Ek}\approx 10^{-3}$ needed to resolve these simulations.